OpenStudy (anonymous):

Find the derivative of the function. MEDAL WILL BE REWARDED!!! f(x)=x(3x-9)^3

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

product rule and chain rule

OpenStudy (anonymous):

okay heres what I got f' (x)=1(3x−9)^3 +81x^3 −486x^2 +729x

hartnn (hartnn):

(3x-9)^3 + 9x (3x-9)^2

hartnn (hartnn):

(3x-9)^2 (3x-9+9x)

OpenStudy (anonymous):

oh okay, so is that my final answer?

hartnn (hartnn):

simplify it...

hartnn (hartnn):

3x + 9x = 12 x

OpenStudy (anonymous):

is there an x on the 12? woops...maybe I added wrong....

hartnn (hartnn):

(3x-9)^2 (3x-9+9x) = (3x-9)^2 (12x-9)

OpenStudy (anonymous):

oh yes your right, sorry wrote it wrong! Thank you so much again!

hartnn (hartnn):

welcome :)